OpenStudy (anonymous):

Solve these two system of equations -3x+7y=6 and 5x-y=10 using linear combinations preferably

OpenStudy (anonymous):

from the second equation y = 5x - 10, inserting this to the first equation: -3x + 7(5x - 10) = 6 -3x+35x -70 = 6 32x =76 x=76/32 = 19/8 Finally insert the value of x in y = 5x - 10

OpenStudy (anonymous):

You will get y = 15/8 thus the answer will be (19/8, 15/8) Do you understand?

OpenStudy (anonymous):

Yes! Thank you very much